Number of workers at the beginning of the year                  8000


Number of workers at the end of the year                        9600


Number of workers left the company during the year               500


Number of workers discharged during the year                    100


Number of workers replaced due to leaving and discharge           700


Additional workers employed for expansion during the year         1500

calculate the labor turnover rate by using separation method, replacement method and flux method,


1
Expert's answer
2021-04-26T19:34:16-0400

Separation method:

Labor turnover rate = 600/((9600 + 8000)/2)×100 = 6.82%.

Replacement method:

Labor turnover rate = 700/(9600 + 8000)/2×100 = 7.95%.

Flux method:

Labor turnover rate = (500 + 700)/((9600 + 8000)/2)×100 = 13.64%.